# Corona rigidity

### v3

@@ 7.2. Quotients of $\mathcal{P}({\mathbb{N}})$ and zero-dimensional remainders
We don’t even know whether Question 7.11 has a positive answer for all ideals 3204-determined by closed approximations, not to mention all $F_{\sigma\delta}$ ideals. Since there has been no progress in this direction since [55], an interested reader is invited to consult [55, §15] for open problems (see however Theorem 7.22).
@@
The following was proved by combining the ideas from [50, §4] and a stratification of the clopen algebra of $\beta X$ with copies of $\mathcal{P}({\mathbb{N}})$.
@@ 7.3.1. The Calkin algebra
We will sketch the proof from [59, §17], based on [58] (see also [60]). Fix an automorphism $\Phi$ of $\mathcal{Q}(H)$. By using the stratification of $\mathcal{Q}(H)$ from §6.4, an initial segment of the proof is concerned with the restrictions of $\Phi$ to reduced products of the form $\mathcal{D}[\mathbf{E}]$. We quickly recap the key ideas from §6.4 while slipping in a few additional definitions. Recall that $\operatorname{Part_{{\mathbb{N}}}}$ is the poset of partitions $\mathbf{E}=\langle E_{j}\colon j\in{\mathbb{N}}\rangle$ of ${\mathbb{N}}$ into finite intervals, ordered by $\mathbf{E}\leq^{*}\mathbf{F}$ if $(\forall^{\infty}n)(\exists m)E_{m}\subseteq F_{n}$. Fixing a basis of $H$ once and for all, to $\mathbf{E}\in\operatorname{Part_{{\mathbb{N}}}}$ one associates a decomposition of $H$ into a direct sum of finite-dimensional subspaces $H=\bigoplus_{n}H_{n}$. Denote the projection to $H_{n}$ by $q^{\mathbf{E}}_{n}$ and let, for $X\subseteq{\mathbb{N}}$, $q_{X}^{\mathbf{E}}=\sum_{n\in X}q_{n}^{\mathbf{E}}$. Let
@@
The implication from (7.3.1) to (7.3.2) uses the methods from §5. In order to use the analog of Lemma 5.8 for finite-dimensional $\mathrm{C}^{*}$-algebras, one needs to replace the matrix algebra components of $\mathcal{D}[\mathbf{E}]$ with finite sets. Identify $\mathcal{D}[\mathbf{E}]$ with $\prod_{n}M_{k(n)}(\mathbb{C})$ (with $k(n)=|E_{n}|$) and choose a finite $2^{-n}$-dense $\mathsf{D}(n)\subseteq M_{k(n)}(\mathbb{C})$ (it is convenient to impose additional properties on $\mathsf{D}_{n}$, see [59, Definition 17.4.2]). Then $\mathsf{D}[\mathbf{E}]=\prod_{n}\mathsf{D}(n)$ is a discretization of $\mathcal{D}[\mathbf{E}]$. Since for $a\in(\mathcal{D}[\mathbf{E}])_{1}$ there is $a^{\prime}\in\mathsf{D}[\mathbf{E}]$ such that $a-a^{\prime}$ is compact, topological triviality on $\mathsf{D}[\mathbf{E}]$ implies topological triviality on $\mathsf{D}[\mathbf{E}]$. By the Ulam-stability (Lemma 5.17 and the analog of Proposition 5.10), (7.3.2) follows. (See [59, §17.4] for the details.)
@@
Thus in order to prove (7.3.2) $\Rightarrow$ (7.3.3) it suffices to start from $\mathbf{E}$ such that $\lim_{n\to\infty}|E_{n}|=\infty$. Moreover, the proof of Lemma 7.14 is flexible enough to transfer other properties of $\Phi$, such as topological triviality, and the existence of an $\varepsilon$-approximation (see below) from $\mathcal{D}[\mathbf{E}]$ to $\mathcal{D}[\mathbf{F}]$ (see [59, Lemma 17.5.3]).
(7.3.3) $\Rightarrow$ (7.3.4) is easy ([60, Proposition 17.5.8]).
Still postponing the technicalities, we handle (7.3.4) $\Rightarrow$ (7.3.5). Let $\mathbf{F}$ be such that each $F_{n}$ is a singleton. Then $A=\mathcal{D}[\mathbf{F}]$ is a maximal abelian $\mathrm{C}^{*}$-subalgebra of $\mathcal{B}(H)$ (this is the atomic masa of $\mathcal{B}(H)$; see [59, §12.3]) isomorphic to $\ell_{\infty}$. Let $u$ be the unitary that implements $\Phi$ on $\mathcal{D}[\mathbf{F}]$. Then $\Phi_{1}=\operatorname{Ad}\pi(u^{*})\circ\Phi$ is an inner automorphism if and only $\Phi$ is, and its restriction to $A$ is equal to the identity. Also, for every $\mathbf{E}$, if $v$ is a unitary that implements $\Phi$ on $\mathcal{F}[\mathbf{E}]$, then $(\operatorname{Ad}u^{*}v)(a)-a$ is compact for all $a\in A$. By the Johnson–Parrot Theorem ([59, Theorem 12.3.2]), there is $v_{0}\in A$ such that $v_{0}-u^{*}v$ is compact. A standard perturbation argument gives a unitary $v_{\mathbf{E}}\in A$ such that $v_{\mathbf{E}}-v_{0}$ is compact.
This shows that, as in Lemma 6.24, $\Phi_{1}$ is induced by a coherent family of unitaries $\mathcal{F}_{\Phi}=\{(\mathbf{E},u_{\mathbf{E}})\}$. The poset $\operatorname{Part_{{\mathbb{N}}}}$ is $\sigma$-directed and every subset of cardinality $\aleph_{1}$ is bounded (this is a consequence of [59, Theorem 9.7.8] and the analogous facts for $({\mathbb{N}}^{\mathbb{N}},\leq^{*})$), and uniformization of coherent families with such an index set is something that $\mathrm{OCA_{T}}$ is really good at. A somewhat elaborate, yet fun, argument shows that $\mathcal{F}_{\Phi}$ is trivial ([59, Theorem 17.8.2]). The trivializing unitary implements $\Phi_{1}$ on $\mathcal{Q}(H)$, and therefore $\Phi$ is inner.
@@
It ‘only’ remains to prove that $\mathrm{OCA_{T}}$ implies (7.3.1). Fix $\mathbf{E}$ such that $\lim_{n}|E_{n}|=\infty$ and a discretization $\mathsf{D}[\mathbf{E}]=\prod_{n}\mathsf{D}(n)$ of $\mathcal{D}[\mathbf{E}]$. Fix $\varepsilon>0$. A function $\theta\colon\mathbf{D}_{X}[\mathbf{E}]\to\mathcal{B}(H)$ is an $\varepsilon$-approximation to $\Phi$ if
@@
$\mathrm{OCA_{\infty}}$ asserts that for every separably metrizable space ${\mathcal{X}}$ and a decreasing sequence of open sets $L^{n}\subseteq[{\mathcal{X}}]^{2}$, for $n\in{\mathbb{N}}$ one of the following alternatives applies.
@@
Moore proved that $\mathrm{OCA_{T}}$ implies $\mathrm{OCA_{\infty}}$ ([114], see also [59, Theorem 8.6.6]). Here is another descendant of Lemma 7.3, via Lemma 7.7 and [50, Proposition 3.12.1].
Lemma 7.16 ([60, Lemma 17.6.3]).
@@
We first fix some notation. Fix a discretization $\mathsf{D}[\mathbf{E}]$. Since $\mathbf{E}$ is fixed, let $q_{Y}=q_{Y}^{\mathbf{E}}$, for $Y\subseteq{\mathbb{N}}$. We also fix a second basis for $H$, and let $r_{n}$ be the canonical projection onto the $n$-th element of the such basis, and $r_{X}$, for $X\subseteq{\mathbb{N}}$, be equal to $\sum_{n\in X}r_{n}$. Further, fix an injection $\chi\colon{\mathbb{N}}\to\{0,1\}^{<{\mathbb{N}}}$ and a lifting $\Phi_{*}$ of $\Phi$ such that $p_{Y}:=\Phi_{*}(q_{Y})$ is a projection for every $Y\subseteq{\mathbb{N}}$ and that $\|\Phi_{*}(a)\|\leq 1$ whenever $\|a\|\leq 1$.
@@
A Kuratowski–Ulam argument and a $\sigma$-Borel $\varepsilon$-approximation on $\mathcal{D}_{X}[\mathbf{E}]$ are used to find a C-measurable $3\varepsilon$-approximation on $\mathcal{D}_{Y}[\mathbf{E}]$ for an infinite $Y\subseteq X$ ([59, Lemma 17.7.2]). A Baire category argument produces an infinite $Z\subseteq Y$ and a continuous $3\varepsilon$-approximation on $\mathcal{D}_{Z}[\mathbf{E}]$. By the analog of the Isometry Trick (Lemma 7.14) $\Phi$ has a continuous $3\varepsilon$-approximation on $\mathcal{D}[\mathbf{E}]$ for all $\varepsilon>0$, and the Jankov–von Neumann Uniformization Theorem can be used to prove that $\Phi$ is topologically trivial on $\mathcal{D}[\mathbf{E}]$ ([59, Lemma 17.4.5]), completing the proof of (7.3.2).
@@
If $\Phi$ sends the atomic masa to a masa unitarily equivalent to it, then $\Phi$ cannot send the unilateral shift to its adjoint (see footnote 44), thus a negative answer implies a negative answer to Question 4.16. However, an adaptation of the construction from [124] may provide a positive answer.
@@ 7.3.2. More coronas
The outline of the proof of Theorem 7.18 follows the proof of Theorem 7.13, but is much more difficult. First, even though one can stratify a general corona $\mathcal{Q}(A)$ (similarly to §6.4), the building blocks consist not of products of matrix algebras, but of products of finite-dimensional Banach spaces instead. The appropriate version of the ‘$\mathrm{OCA_{T}}$ lifting theorem’ for reduced products of finite-dimensional spaces appears in [111, Lemma 5.10]. Secondly, the assumptions of the isometry trick do not apply in most coronas (Lemma 7.14). The latter obstruction is the reason why we need to assume Martin’s Axiom in Theorem 7.18.
@@
7.4. Rigidity for fields, linear orderings, trees, and sufficiently random graphs
In Theorem 6.15 we saw that a dividing line for the first part of Conjecture 2.6 is the stability of the theory of the reduced product. We do not have a clean dividing line for the second part of this conjecture, but it has been confirmed in several categories. So far we have seen that the second part of this conjecture holds for quotient Boolean algebras $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$ for certain analytic ideals ${\mathcal{I}}$ as well as for coronas of separable $\mathrm{C}^{*}$-algebras. The case of semilattices $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$ was considered in [94], [90], and [52]. Rigidity results for other categories of discrete axiomatizable structures appear in [32]. The following example is very suggestive. If $F=\prod_{n}F_{n}/{\mathcal{I}}$ is a reduced product of fields, then the idempotents in $F$ form a copy of $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$, with the multiplication and addition (on disjoint sets) interpreted as the Boolean algebraic intersection and union. Thus we have a definable copy of the Boolean algebra $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$, and to an isomorphism $\Phi$ between such reduced products one can associate an isomorphism $\alpha_{\Phi}$ between quotients of the form $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$. Triviality of the latter is the first step in a proof (using $\mathrm{OCA_{T}}$ and $\mathrm{MA}$) that $\Phi$ is (with an appropriate definition) trivial. A localization of this argument gives rigidity in reduced products of structures in some other categories (linear orderings, trees, sufficiently random graphs) in which there is no obvious ‘global’ copy of $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$. These rigidity results came as a surprise (even to the authors) and they also provided a slightly more optimal proof that all automorphisms of $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ are trivial (see ‘The role of MA’ in §11.0.2). The continuous version of the main rigidity result of [32] appears in [34].

### v4

@@ 7.2. Quotients of $\mathcal{P}({\mathbb{N}})$ and zero-dimensional remainders
We don’t even know whether Question 7.11 has a positive answer for all $F_{\sigma\delta}$ ideals (see however Theorem 7.22). After two decades with virtually no progress, in [64] the conclusions of Theorem 7.8 and Corollary 7.9 have been strengthened to include all ideals that satisfy the property from Question 7.10 with $d=80$ and their assumptions have been weakened to $\mathrm{OCA_{T}}$ and $\mathrm{MA_{\sigma-linked}}$ (see §11.0.2).
@@
The following was proved by combining the ideas from [52, §4] and a stratification of the clopen algebra of $\beta X$ with copies of $\mathcal{P}({\mathbb{N}})$.
@@ 7.3.1. The Calkin algebra
We will sketch the proof from [61, §17], based on [60] (see also [62]). Fix an automorphism $\Phi$ of $\mathcal{Q}(H)$. By using the stratification of $\mathcal{Q}(H)$ from §6.4, an initial segment of the proof is concerned with the restrictions of $\Phi$ to reduced products of the form $\mathcal{D}[\mathbf{E}]$. We quickly recap the key ideas from §6.4 while slipping in a few additional definitions. Recall that $\operatorname{Part_{{\mathbb{N}}}}$ is the poset of partitions $\mathbf{E}=\langle E_{j}\colon j\in{\mathbb{N}}\rangle$ of ${\mathbb{N}}$ into finite intervals, ordered by $\mathbf{E}\leq^{*}\mathbf{F}$ if $(\forall^{\infty}n)(\exists m)E_{m}\subseteq F_{n}$. Fixing a basis of $H$ once and for all, to $\mathbf{E}\in\operatorname{Part_{{\mathbb{N}}}}$ one associates a decomposition of $H$ into a direct sum of finite-dimensional subspaces $H=\bigoplus_{n}H_{n}$. Denote the projection to $H_{n}$ by $q^{\mathbf{E}}_{n}$ and let, for $X\subseteq{\mathbb{N}}$, $q_{X}^{\mathbf{E}}=\sum_{n\in X}q_{n}^{\mathbf{E}}$. Let
@@
The implication from (7.3.1) to (7.3.2) uses the methods from §5. In order to use the analog of Lemma 5.8 for finite-dimensional $\mathrm{C}^{*}$-algebras, one needs to replace the matrix algebra components of $\mathcal{D}[\mathbf{E}]$ with finite sets. Identify $\mathcal{D}[\mathbf{E}]$ with $\prod_{n}M_{k(n)}(\mathbb{C})$ (with $k(n)=|E_{n}|$) and choose a finite $2^{-n}$-dense $\mathsf{D}(n)\subseteq M_{k(n)}(\mathbb{C})$ (it is convenient to impose additional properties on $\mathsf{D}_{n}$, see [61, Definition 17.4.2]). Then $\mathsf{D}[\mathbf{E}]=\prod_{n}\mathsf{D}(n)$ is a discretisation of $\mathcal{D}[\mathbf{E}]$. Since for $a\in(\mathcal{D}[\mathbf{E}])_{1}$ there is $a^{\prime}\in\mathsf{D}[\mathbf{E}]$ such that $a-a^{\prime}$ is compact, topological triviality on $\mathsf{D}[\mathbf{E}]$ implies topological triviality on $\mathsf{D}[\mathbf{E}]$. By the Ulam-stability (Lemma 5.17 and the analog of Proposition 5.10), (7.3.2) follows. (See [61, §17.4] for the details.)
@@
Thus in order to prove (7.3.2) $\Rightarrow$ (7.3.3) it suffices to start from $\mathbf{E}$ such that $\lim_{n\to\infty}|E_{n}|=\infty$. Moreover, the proof of Lemma 7.14 is flexible enough to transfer other properties of $\Phi$, such as topological triviality, and the existence of an $\varepsilon$-approximation (see below) from $\mathcal{D}[\mathbf{E}]$ to $\mathcal{D}[\mathbf{F}]$ (see [61, Lemma 17.5.3]).
(7.3.3) $\Rightarrow$ (7.3.4) is easy ([62, Proposition 17.5.8]).
Still postponing the technicalities, we handle (7.3.4) $\Rightarrow$ (7.3.5). Let $\mathbf{F}$ be such that each $F_{n}$ is a singleton. Then $A=\mathcal{D}[\mathbf{F}]$ is a maximal abelian $\mathrm{C}^{*}$-subalgebra of $\mathcal{B}(H)$ (this is the atomic masa of $\mathcal{B}(H)$; see [61, §12.3]) isomorphic to $\ell_{\infty}$. Let $u$ be the unitary that implements $\Phi$ on $\mathcal{D}[\mathbf{F}]$. Then $\Phi_{1}=\operatorname{Ad}\pi(u^{*})\circ\Phi$ is an inner automorphism if and only $\Phi$ is, and its restriction to $A$ is equal to the identity. Also, for every $\mathbf{E}$, if $v$ is a unitary that implements $\Phi$ on $\mathcal{F}[\mathbf{E}]$, then $(\operatorname{Ad}u^{*}v)(a)-a$ is compact for all $a\in A$. By the Johnson–Parrot Theorem ([61, Theorem 12.3.2]), there is $v_{0}\in A$ such that $v_{0}-u^{*}v$ is compact. A standard perturbation argument gives a unitary $v_{\mathbf{E}}\in A$ such that $v_{\mathbf{E}}-v_{0}$ is compact.
This shows that, as in Lemma 6.24, $\Phi_{1}$ is induced by a coherent family of unitaries $\mathcal{F}_{\Phi}=\{(\mathbf{E},u_{\mathbf{E}})\}$. The poset $\operatorname{Part_{{\mathbb{N}}}}$ is $\sigma$-directed and every subset of cardinality $\aleph_{1}$ is bounded (this is a consequence of [61, Theorem 9.7.8] and the analogous facts for $({\mathbb{N}}^{\mathbb{N}},\leq^{*})$), and uniformisation of coherent families with such an index set is something that $\mathrm{OCA_{T}}$ is really good at. A somewhat elaborate, yet fun, argument shows that $\mathcal{F}_{\Phi}$ is trivial ([61, Theorem 17.8.2]). The trivializing unitary implements $\Phi_{1}$ on $\mathcal{Q}(H)$, and therefore $\Phi$ is inner.
@@
It ‘only’ remains to prove that $\mathrm{OCA_{T}}$ implies (7.3.1). Fix $\mathbf{E}$ such that $\lim_{n}|E_{n}|=\infty$ and a discretisation $\mathsf{D}[\mathbf{E}]=\prod_{n}\mathsf{D}(n)$ of $\mathcal{D}[\mathbf{E}]$. Fix $\varepsilon>0$. A function $\theta\colon\mathbf{D}_{X}[\mathbf{E}]\to\mathcal{B}(H)$ is an $\varepsilon$-approximation to $\Phi$ if
@@
$\mathrm{OCA_{\infty}}$ asserts that for every separably metrisable space ${\mathcal{X}}$ and a decreasing sequence of open sets $L^{n}\subseteq[{\mathcal{X}}]^{2}$, for $n\in{\mathbb{N}}$ one of the following alternatives applies.
@@
Moore proved that $\mathrm{OCA_{T}}$ implies $\mathrm{OCA_{\infty}}$ ([120], see also [61, Theorem 8.6.6]). Here is another descendant of Lemma 7.3, via Lemma 7.7 and [52, Proposition 3.12.1].
Lemma 7.16 ([62, Lemma 17.6.3]).
@@
We first fix some notation. Fix a discretisation $\mathsf{D}[\mathbf{E}]$. Since $\mathbf{E}$ is fixed, let $q_{Y}=q_{Y}^{\mathbf{E}}$, for $Y\subseteq{\mathbb{N}}$. We also fix a second basis for $H$, and let $r_{n}$ be the canonical projection onto the $n$-th element of the such basis, and $r_{X}$, for $X\subseteq{\mathbb{N}}$, be equal to $\sum_{n\in X}r_{n}$. Further, fix an injection $\chi\colon{\mathbb{N}}\to\{0,1\}^{<{\mathbb{N}}}$ and a lifting $\Phi_{*}$ of $\Phi$ such that $p_{Y}:=\Phi_{*}(q_{Y})$ is a projection for every $Y\subseteq{\mathbb{N}}$ and that $\|\Phi_{*}(a)\|\leq 1$ whenever $\|a\|\leq 1$.
@@
A Kuratowski–Ulam argument and a $\sigma$-Borel $\varepsilon$-approximation on $\mathcal{D}_{X}[\mathbf{E}]$ are used to find a C-measurable $3\varepsilon$-approximation on $\mathcal{D}_{Y}[\mathbf{E}]$ for an infinite $Y\subseteq X$ ([61, Lemma 17.7.2]). A Baire category argument produces an infinite $Z\subseteq Y$ and a continuous $3\varepsilon$-approximation on $\mathcal{D}_{Z}[\mathbf{E}]$. By the analog of the Isometry Trick (Lemma 7.14) $\Phi$ has a continuous $3\varepsilon$-approximation on $\mathcal{D}[\mathbf{E}]$ for all $\varepsilon>0$, and the Jankov–von Neumann Uniformisation Theorem can be used to prove that $\Phi$ is topologically trivial on $\mathcal{D}[\mathbf{E}]$ ([61, Lemma 17.4.5]), completing the proof of (7.3.2).
@@
If $\Phi$ sends the atomic masa to a masa unitarily equivalent to it, then $\Phi$ cannot send the unilateral shift to its adjoint (see footnote 45), thus a negative answer implies a negative answer to Question 4.16. However, an adaptation of the construction from [130] may provide a positive answer.
@@ 7.3.2. More coronas
The outline of the proof of Theorem 7.18 follows the proof of Theorem 7.13, but is much more difficult. First, even though one can stratify a general corona $\mathcal{Q}(A)$ (similarly to §6.4), the building blocks consist not of products of matrix algebras, but of products of finite-dimensional Banach spaces instead. The appropriate version of the ‘$\mathrm{OCA_{T}}$ lifting theorem’ for reduced products of finite-dimensional spaces appears in [117, Lemma 5.10]. Secondly, the assumptions of the isometry trick do not apply in most coronas (Lemma 7.14). The latter obstruction is the reason why we need to assume Martin’s Axiom in Theorem 7.18.
@@
7.4. Rigidity in other categories
In Theorem 6.15 we saw that a dividing line for the first part of Conjecture 2.6 is the stability of the theory of the reduced product. We do not have a clean dividing line for the second part of this conjecture, but it has been confirmed in several categories. So far we have seen that the second part of this conjecture holds for quotient Boolean algebras $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$ for certain analytic ideals ${\mathcal{I}}$ as well as for coronas of separable $\mathrm{C}^{*}$-algebras. The case of semilattices $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$ was considered in [98], [94], and [54]. Rigidity results for some other categories of discrete structures appear in [32]. The following example, suggested to us by Zoé Chatzidakis, is suggestive. If $F=\prod_{n}F_{n}/{\mathcal{I}}$ is a reduced product of fields, then the idempotents in $F$ form a copy of $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$, with the multiplication and addition (on disjoint sets) interpreted as the Boolean algebraic intersection and union. Thus we have a definable copy of the Boolean algebra $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$, and to an isomorphism $\Phi$ between such reduced products one can associate an isomorphism $\alpha_{\Phi}$ between quotients of the form $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$. Triviality of the latter is the first step in a proof (using $\mathrm{OCA_{T}}$ and $\mathrm{MA}$) that $\Phi$ is (with a natural definition) trivial. This observation was formalised as ‘theory recognizing coordinates’ in [32, Definition 2.6], giving rigidity in reduced products of structures in some other categories (linear orderings, trees, sufficiently random graphs) in which there is no obvious ‘global’ copy of $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$. These rigidity results came as a surprise (even to the authors) and they also led to more optimal proofs that all automorphisms of $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ are trivial and of the OCA Lifting Theorem (see §11.0.2). In [66, Theorem 1] it was proved that a category (not necessarily axiomatisable) of countable structures recognises coordinates (and therefore $\mathrm{OCA_{T}}$ and $\mathrm{MA}$ imply that all automorphisms of reduced products over $\operatorname{Fin}$ of its elements are trivial) if and only if the formula $x=x^{\prime}\rightarrow y=y^{\prime}$ is equivalent to an $h$-formula (see §6.1.6). One consequence is that a category recognises coordinates if and only if its theory does. This paper, true to its title, contains other surprises on the theory of recognizing coordinates.
The continuous version of the main rigidity result of [32] appears in [35] and it was applied to universal sofic groups and Higson coronas, respectively, in [33] and [176].